OpenStudy (anonymous):

Which factor do you cancel from the numerator and denominator to simplify x^2 - 3x - 4 / x^2 - 16?

OpenStudy (anonymous):

Is the answer x - 4?

OpenStudy (anonymous):

\[\frac{ x^2-3x-4 }{ x^2-16 }\] then factor both the top and the bottom \[\frac{ (x-4)(x+1) }{ (x-4)(x+4) }\] \[\frac{ (x-4) }{ (x-4) }\] cancels out and you get\[\frac{ (x+1) }{ (x+4) }\] so yes (x-4) is canceled out